1. A small electric immersion heater is used to heat 68 g of water for a cup of instant coffee. The heater is labeled “76 watts” (it converts electrical energy to thermal energy at this rate). Calculate the time required to bring all this water from 23°C to 100°C, ignoring any heat losses. (The specific heat of water is 4186 J/kg·K.)


  1. A vertical glass tube of length L = 1.3m is half filled with a liquid at 15.8 ˚C. How much will the height of the liquid column change when the tube is heated to 34.0 ˚C? Take αglass = 1.25× 10-5/K and βliquid = 4.46× 10-5/K.
1
Expert's answer
2022-01-30T13:52:35-0500

For the first problem:

"t=\dfrac{cm\Delta t}{P}=288\text{ s, or 4 min 48 s.}"


For another problem:


"\Delta h=\dfrac L2\Delta t(\alpha+\beta)=3.37·10^{-4}\text{ m}."